Please help with the following...Thanks
Find the point-slope form of the line satisfying the given conditions and use this to write the slope-intercept form of the equation.
Passing through (-3, -4), (-4, -3)

Answer by checkley75(3666)

SLOPE=(Y2-Y1)/(X2-X1)=(-3+4)/(-4+3)=1/-1=-1 FOR THE SLOPE (m).
NOW REPLACE THE X & Y VALUES WITH ONE SET OF X,Y POINTS & SOLVE FOR THE Y INTERCEPT(b)
Y=mX+b
-4=-1*-3+b
b=-4-3
b=-7 SO THE LINE EQUATION IS;
Y=-X-7